# Occupied Processes: Going with the Flow

### v3

@@ Abstract
A stochastic process $X$ becomes occupied when it is enlarged with its occupation flow $\operatorname{\mathcal{O}}$, which tracks the time spent by the path at each level. Crucially, the occupied process $(\operatorname{\mathcal{O}},X)$ enjoys a Markov structure when $X$ is Markov. We develop a novel Itô calculus for occupied processes that lies midway between Dupire’s functional Itô calculus and the classical version. We derive a surprisingly simple Itô formula and, through Feynman-Kac, unveil a broad class of path-dependent PDEs where $\operatorname{\mathcal{O}}$ plays the role of time. The space variable, given by the current value of $X$, remains finite-dimensional, thereby paving the way for standard elliptic PDE techniques and numerical methods.
In the financial applications, we demonstrate that occupation flows provide unified Markovian lifts for exotic options and variance instruments, allowing financial institutions to price and manage derivatives books with a single numerical solver. We then explore avenues in financial modeling where volatility is driven by the occupied process. In particular, we propose the local occupied volatility (LOV) model which not only calibrates to European vanilla options but also offers the flexibility to capture stylized facts of volatility and fit other instruments such as VIX options or American put options on non-dividend-paying single names. We also present an extension of forward variance models that leverages the entire forward occupation surface.
@@ 3.4 Occupied SDEs and Dirichlet Problems
Incorporating the occupation flow into stochastic differential equations is uncommon in the literature. A notable exception is the class of self-attracting diffusions introduced by Cranston and Le Jan [29], inspired by the Brownian polymer model of Durrett and Rogers [41]. Therein, the drift is influenced by the the calendar time occupation flow so as to create ergodic behaviors. These diffusions were further studied by Raimond [72], and extended in [6, 7, 8, 9] to self-interacting diffusions, also including SDEs exhibiting self-repulsion. Whereas self-interacting diffusions use the occupation flow in the drift, note that OSDEs involve the occupation in the diffusion coefficient as well. Besides, let us mention the recent papers [23, 37] linking self-interacting diffusions to ergodic McKean-Vlasov SDEs.
@@ 4.1 Exotic Options and Structured Products
assuming the law $\operatorname{\mathbb{Q}}\circ X_{T}^{-1}$ has no atoms. However, only finitely many strikes and maturities are available in practice. Even when the range accrual is monitored daily, so the right side of (52) becomes a sum, options on the reference rate rarely expires on every business day. Hence, the valuation of range accruals would still benefit from the unified pricing and modeling framework in Fig. 3.
@@
Parisian options [24, 62] were introduced to relax the knock-in/out component of barrier options seen in Example 8. Precisely, the option is activated or terminated only once the asset has performed an excursion above/below the barrier of length at least $\delta>0$, referred to as the window. A variant, called cumulative Parisian or Parasian option [57], requires that the cumulative time above/below the barrier exceeds the window. As the payoff of excursion-based Parisian options depends on the chronology of the path, the occupied lift can only be applied to the cumulative counterparts in view of Theorem A.1. For example, consider a cumulative, asset-or-nothing, up-and-out Parisian call option with strike $K$, barrier $B>K$, and window $\delta>0$. Then its payoff can be rewritten as
$\xi=\begin{cases}X_{T},&\text{ if }\int_{0}^{T}\mathds{1}_{\{X_{t}\geq B\}}dt<\delta,\\ 0,&\text{otherwise,}\end{cases}=X_{T}\mathds{1}_{\{\tilde{\operatorname{\mathcal{O}}}_{T}(A)<\delta\}}=\phi(\tilde{\operatorname{\mathcal{O}}}_{T}(A),X_{T}),\quad A=[B,\infty).$
The introduction of a window is particularly appealing to the buyer of knock-out Parisian options as it (partially) prevents eventual price manipulation from the seller to push the underlying towards the barrier thus making the contracts worthless. From the seller’s perspective, Parisian options are easier to hedge as the soft barrier effectively smoothen the Greeks. Other instruments exhibit a similar relaxation feature. One example is the soft call provision of convertible bonds, where the issuer discontinues (calls) the contract once the underlying has spent a prescribed amount of time above a threshold.
@@ 5 Occupied Volatility
$\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma(\tilde{\operatorname{\mathcal{O}}}_{t},X_{t})^{2}\ |\ X_{t}]=\sigma_{\text{loc}}(t,X_{t})^{2}.$ (62)
@@ 5.2 Local Occupied Volatility (LOV) Model
for some path-dependent feature $F_{t}(\omega)=\textnormal{f}(t,\omega)\in\operatorname{\mathfrak{F}}$, $\textnormal{f}\mathrel{\mathop{\ordinarycolon}}\Omega_{T}\to\operatorname{\mathfrak{F}}$. The feature space $\operatorname{\mathfrak{F}}$ may be finite-dimensional, e.g., $F_{t}\in\{\sup_{s\leq t}X_{s},\int_{0}^{t}X_{s}ds\}$ in Brunick and Shreve [20],or infinite-dimensional in occupied volatility models where $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$; see Section 5.2.1. Recall the smile calibration condition
@@
and introduce the projected feature $\hat{F}_{t}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[F_{t}|X_{t}]=\hat{f}(t,X_{t})$ for some map $\hat{f}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}_{+}^{2}\to\operatorname{\mathfrak{F}}.$ If $\sigma^{2}(F,X)$ were linear in $F$, then $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(F_{t},X_{t})|X_{t}]=\sigma^{2}(\hat{F}_{t},X_{t})$, and (69) would imply that $\sigma_{\text{loc}}(t,X_{t})=\sigma(\hat{f}(t,X_{t}),X_{t})$. In general, one would linearize $\sigma^{2}$ using a first order Markovian expansion around $\hat{F}_{t}$, namely
@@
with the Fréchet derivative $D\sigma^{2}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathfrak{F}}\to\operatorname{\mathfrak{F}}^{*}$ and dual pairing $(\ell,F)\in\operatorname{\mathfrak{F}}^{*}\times\operatorname{\mathfrak{F}}\ni(\ell,F)\to\ell\cdot F$, provided that $\operatorname{\mathfrak{F}}$ is a linear space. The derivative $D\sigma^{2}(\hat{F}_{t},X_{t})$ can be seen as the sensitivity of variance with respect to path-dependent shocks. Noting that $\sigma^{2}(\hat{F}_{t},X_{t})^{2}$, $D\sigma^{2}(\hat{F}_{t},X_{t})$ are functions of $(t,X_{t})$, we can postulate that
@@
where $\sigma^{2}(\hat{F}_{t},X_{t})$ in (70) is replaced by the local variance. Since $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[F_{t}-\hat{F}_{t}|X_{t}]=0$, the above specification automatically satisfies the smile calibration condition (69) for any choice of $\ell$ such that the instantaneous variance remains positive. The local path-dependent volatility (LPDV) model (71) is reminiscent of local stochastic volatility (LSV) models [54, 55] where one multiplies $\sigma_{\text{loc}}$ by an exogenous stochastic processes $\alpha_{t}$ and normalized such that (69) holds. The volatility process is thus of the form
@@
where the conditional expectation is typically estimated using the particle method of Guyon and Henry-Labordère [54]. It is a powerful family of models that is consistent with vanilla options and capable of producing more realistic spot-vol dynamics than the standard local volatility model. On the other hand, LPDV models also preserve market completeness, and explains the dynamics of volatility directly from the price history. Moreover, considering an additive correction term $\ell(t,X_{t})\cdot(F_{t}-\hat{F}_{t})$ term is arguably preferred numerically over the fraction in (72) with distribution-dependent denominator.
@@ 5.2.1 The LOV Model
Note that $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\frac{d}{dh}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}+h\delta_{x},X_{t})|_{h=0}$ describes the change in instantaneous variance when the asset spends extra time at some level $x$. We then replace $\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})$ in (73) by the local variance and write $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\gamma_{t}^{\kappa}\ell(t,X_{t},x)$ with the inverse total mass $\gamma_{t}^{\kappa}=\operatorname{\mathcal{O}}_{t}^{\kappa}(\operatorname{\mathbb{R}})^{-1}=\frac{\kappa}{e^{\kappa t}-1}$ if $\kappa>0$ and $\gamma_{t}^{0}=\frac{1}{t}$. This leads finally to the Local Occupied Volatility (LOV) model,
@@
The factor $\gamma_{t}^{\kappa}$ is introduced so that the magnitude of $\ell$ remains constant over time, and allows to interpreted the linear term as an average sensitivity of instantaneous variance as the realized occupation measure $\mathcal{O}_{t}^{\kappa}$ deviates from the implied one $\hat{\mathcal{O}}_{t}^{\kappa}$. The LOV model may be further simplified by assuming that $\ell$ is time-independent and, aligned with our experiments with neural OSDEs in the next section, that it depends on $(X_{t},x)$ only through the ratio $x/X_{t}$; see Example 14.
@@ 5.3 Neural Calibration and Occupation Sensitivities
where $\partial_{m}\sigma^{\theta}$ denotes the partial derivative with respect to $O_{m}$. Some illustrations of $x\mapsto\ell(t,X_{t},x)$ will be given in Section 5.3.3.
@@ 5.3.3 Path Dependence and Occupation Sensitivities
Fig. 11 Displays the volatility values $\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ against the spot $X_{t}^{\theta}$ for $2^{10}$ trajectories on the full calibration horizon ($t\approx 0.5$). Observe that the dots do not lie on the graph of $x\to\hat{\sigma}^{\theta}(t,x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa,\theta},X^{\theta})|=x]^{1/2}$, confirming the presence of path dependence in $\sigma^{\theta}$. Moreover, the volatility level typically increases as $X_{t}^{\theta}$, in line with the leverage effect. Note also that $\hat{\sigma}^{\theta}(t,x)$ is a smooth approximation of the local volatility function since the neural occupation volatility has been calibrated to the vanilla options.
Figure 11: Neural volatility $\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ versus the spot $X_{t}^{\theta}$ for $2^{10}$ trajectories.
@@ 6 Forward Occupation Models
where $\text{OTM}(K,T)$ is the price of the corresponding out-of-the-money vanilla contract (put if $K<S_{t}$ and call otherwise). As $\mathfrak{o}_{0}^{T}=\partial_{T}\operatorname{\mathcal{O}}_{0}^{T}$, the instantaneous forward occupation surface can be extracted from the weighted sum of calendar spreads $\mathfrak{o}_{0}^{T}(A)=2\int_{A}\partial_{T}\text{OTM}(K,T)\frac{dK}{K^{2}}.$ An illustration is given in Fig. 13 for SPX options. In conclusion, one can can add granularity to forward variance models by including all the forward occupation times $\operatorname{\mathcal{O}}_{o}^{\cdot}(A)$, which are tradeable quantities as well.
@@ 6.1 Dynamics of Forward Occupation Flows
$\int_{0}^{t}\eta^{T}_{u}(A\cup A^{\prime})\cdot dW_{u}=\mathfrak{o}_{t}^{T}(A\cup A^{\prime})=\mathfrak{o}_{t}^{T}(A)+\mathfrak{o}_{t}^{T}(A^{\prime})=\int_{0}^{t}(\eta^{T}_{u}(A)+\eta^{T}_{u}(A^{\prime}))\cdot dW_{u}.$
@@
with volatility $H_{t}^{T}=\int_{t}^{T}\eta_{t}^{s}ds=\int_{t}^{T}\eta(t,s,\mathfrak{o}_{t}^{\tau(t,s)})ds\in\operatorname{\mathcal{M}}^{d}$. Unlike $\operatorname{\mathcal{O}}$, the forward occupation flow $\operatorname{\mathcal{O}}^{T}$ does not play the role of time since its total mass process, the variance swap rate $\text{VS}^{T}$, is of infinite variation. Moreover, by leveraging the tools in Section 3, we obtain the following chain rule for functionals of $\operatorname{\mathcal{O}}^{T}$, namely
@@
$\varphi(\operatorname{\mathcal{O}}_{T})=(\operatorname{\mathcal{O}}_{T}(A^{\pm})-K)^{+},\quad A^{-}=(-\infty,X_{0}],\;A^{+}=(X_{0},\infty).$

### v4

@@ Abstract
A stochastic process $X$ becomes occupied when it is enlarged with its occupation flow $\mathcal{O}$, which tracks the time spent by the path at each level. Crucially, the occupied process $(\mathcal{O},X)$ enjoys a Markov structure when $X$ is Markov. We develop a novel Itô calculus for occupied processes that lies midway between Dupire’s functional Itô calculus and the classical version. We derive a surprisingly simple Itô formula and, through Feynman-Kac, unveil a broad class of path-dependent PDEs where $\mathcal{O}$ plays the role of time. The space variable, given by the current value of $X$, remains finite-dimensional, thereby paving the way for standard elliptic PDE techniques and numerical methods.
In the financial applications, we demonstrate that occupation flows provide unified Markovian lifts for exotic options and variance instruments, allowing financial institutions to price and manage derivatives books with a single numerical solver. We then explore avenues in financial modeling where volatility is driven by the occupied process. In particular, we propose the local occupied volatility (LOV) model which not only calibrates to European vanilla options but also offers the flexibility to capture stylized facts of volatility and fit other instruments. We also present an extension of forward variance models that leverages the entire forward occupation surface.
@@ 3.4 Occupied SDEs and Dirichlet Problems
Incorporating the occupation flow into stochastic differential equations is uncommon in the literature. A notable exception is the class of self-attracting diffusions of Cranston and Le Jan [29], inspired by the Brownian polymer model of Durrett and Rogers [41]. Therein, the drift is influenced by the the calendar time occupation flow so as to create ergodic behaviors. These diffusions were further studied by Raimond [72], and extended in [6, 7, 8, 9] to self-interacting diffusions, also including SDEs exhibiting self-repulsion. Whereas self-interacting diffusions use the occupation flow in the drift, note that OSDEs involve the occupation in the diffusion coefficient as well. Besides, let us mention the recent papers [23, 37] linking self-interacting diffusions to ergodic McKean-Vlasov SDEs.
@@ 4.1 Exotic Options and Structured Products
assuming that $\operatorname{\mathbb{Q}}\circ X_{T}^{-1}$ has no atoms. However, only finitely many strikes and maturities are available in practice. Even when the range accrual is monitored daily, so the right side of (52) becomes a sum, options on the reference rate rarely expires on every business day. Hence, the valuation of range accruals would still benefit from the unified pricing and modeling framework in Fig. 3.
@@
Parisian options [24, 62] were introduced to relax the knock-in/out component of barrier options seen in Example 8. Precisely, the option is activated or terminated only once the asset has performed an excursion above/below the barrier of length at least $\delta>0$, referred to as the window. A variant, called cumulative Parisian or Parasian option [57], requires that the cumulative time above/below the barrier exceeds the window. As the payoff of excursion-based Parisian options depends on the chronology of the path, the occupied lift can only be applied to the cumulative counterparts in view of Theorem A.1. For example, consider an asset-or-nothing, up-and-out cumulative Parisian call option with strike $K$, barrier $B>K$, and window $\delta>0$, namely
$\xi=\begin{cases}X_{T},&\text{ if }\int_{0}^{T}\mathds{1}_{\{X_{t}\geq B\}}dt<\delta,\\ 0,&\text{otherwise.}\end{cases}$
Letting $A=[B,\infty)$, we can rewrite the above payoff as $\xi=X_{T}\mathds{1}_{\{\tilde{\operatorname{\mathcal{O}}}_{T}(A)<\delta\}},$ which is indeed of the form (51). The relaxed knock-in/out feature is particularly appealing to the buyer of knock-out Parisian options as it (partially) prevents price manipulation from the seller to push the underlying towards the barrier, making the contracts worthless. From the seller’s perspective, Parisian options are easier to hedge as the soft barrier effectively smoothen the Greeks. Other instruments exhibit a similar relaxation feature. One example is the soft call provision of convertible bonds, where the issuer discontinues (calls) the contract once the underlying has spent a prescribed amount of time above a threshold.
@@ 5 Occupied Volatility
$\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(\tilde{\operatorname{\mathcal{O}}}_{t},X_{t})\ |\ X_{t}]=\sigma_{\text{loc}}^{2}(t,X_{t}).$ (62)
@@ 5.2 Local Occupied Volatility (LOV) Model
for some path-dependent feature $F_{t}(\omega)=\textnormal{f}(t,\omega)\in\operatorname{\mathfrak{F}}$, $\textnormal{f}\mathrel{\mathop{\ordinarycolon}}[0,T]\times\Omega_{T}\to\operatorname{\mathfrak{F}}$. The feature space $\operatorname{\mathfrak{F}}$ may be finite-dimensional, e.g., $F_{t}\in\{\sup_{s\leq t}X_{s},\int_{0}^{t}X_{s}ds\}$ in Brunick and Shreve [20],or infinite-dimensional in occupied volatility models where $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$; see Section 5.2.1. Recall the smile calibration condition
@@
and introduce the projected feature $\hat{F}_{t}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[F_{t}|X_{t}]=\hat{f}(t,X_{t})$ for some map $\hat{f}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}_{+}^{2}\to\operatorname{\mathfrak{F}}.$ If $\sigma^{2}(F,X)$ were linear in $F$, then $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(F_{t},X_{t})|X_{t}]=\sigma^{2}(\hat{F}_{t},X_{t})$, and (69) would imply that $\sigma_{\text{loc}}(t,X_{t})=\sigma(\hat{f}(t,X_{t}),X_{t})$. In general, one would linearize $F\mapsto\sigma^{2}(F,X)$ using a first order Markovian expansion around $\hat{F}_{t}$, namely
@@
with the Fréchet derivative $D\sigma^{2}(\cdot,X_{t})\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathfrak{F}}\to\operatorname{\mathfrak{F}}^{*}$ and dual pairing $\operatorname{\mathfrak{F}}^{*}\times\operatorname{\mathfrak{F}}\ni(\ell,F)\to\ell\cdot F$, provided that $\operatorname{\mathfrak{F}}$ is a linear space. The derivative $D\sigma^{2}(\hat{F}_{t},X_{t})$ can be seen as the sensitivity of variance with respect to non-Markovian shocks. Noting that $\sigma^{2}(\hat{F}_{t},X_{t})^{2}$, $D\sigma^{2}(\hat{F}_{t},X_{t})$ are functions of $(t,X_{t})$, we can postulate that
@@
by replacing $\sigma^{2}(\hat{F}_{t},X_{t})$ in (70) with the local variance. Since $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[F_{t}-\hat{F}_{t}|X_{t}]=0$, the above specification automatically satisfies the smile calibration condition (69) for any reasonable choice of $\ell$, e.g., such that $\sigma^{2}(F_{t},X_{t})$ remains positive. The local path-dependent volatility model (71) is reminiscent of local stochastic volatility (LSV) models [54, 55] where one multiplies $\sigma_{\text{loc}}$ by an exogenous stochastic processes $\alpha_{t}$ and normalized such that (69) holds. The volatility process is thus of the form
@@
where the conditional expectation is typically estimated using the particle method of Guyon and Henry-Labordère [54]. It is a powerful family of models that is consistent with vanilla options and capable of producing more realistic spot-vol dynamics than the standard local volatility model. On the other hand, local path-dependent volatility models also preserve market completeness by explaining future volatility from the price history directly. Moreover, considering an additive correction term $\ell(t,X_{t})\cdot(F_{t}-\hat{F}_{t})$ term is arguably preferred numerically over the fraction in (72) with distribution-dependent denominator.
@@ 5.2.1 The LOV Model
Note that $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\frac{d}{dh}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}+h\delta_{x},X_{t})|_{h=0}$ describes the change in instantaneous variance when the asset spends extra time at some level $x$. We then replace $\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})$ in (73) by the local variance and write $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\gamma_{t}^{\kappa}\ell(t,X_{t},x)$, with the inverse total mass $\gamma_{t}^{\kappa}=\operatorname{\mathcal{O}}_{t}^{\kappa}(\operatorname{\mathbb{R}})^{-1}=\frac{\kappa}{e^{\kappa t}-1}$ if $\kappa>0$ and $\gamma_{t}^{0}=\frac{1}{t}$. This leads finally to the Local Occupied Volatility (LOV) model,
@@
The factor $\gamma_{t}^{\kappa}$ is introduced so that the magnitude of $\ell$ remains constant over time, and allows to interpreted the linear term as an average sensitivity of instantaneous variance as the realized occupation measure $\mathcal{O}_{t}^{\kappa}$ deviates from the implied one $\hat{\mathcal{O}}_{t}^{\kappa}$. The LOV model may be further simplified by assuming that $\ell$ is time-independent and, aligned with our experiments with neural OSDEs in the next section, that it depends on $X_{t},x$ only through the ratio $x/X_{t}$; see Example 14.
@@ 5.3 Neural Calibration and Occupation Sensitivities
where $\partial_{m}$ denotes the partial derivative with respect to $O_{m}$. Illustrations of $x\mapsto\ell(t,X_{t},x)$ will be given in Section 5.3.3.
@@ 5.3.3 Path Dependence and Occupation Sensitivities
Fig. 11 Displays the volatility values $\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ against the spot $X_{t}^{\theta}$ for $2^{10}$ trajectories on the full calibration horizon ($t\approx 0.5$). Observe that the dots do not lie on the graph of $\hat{\sigma}^{\theta}(t,\cdot)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa,\theta},X^{\theta}_{t})|X_{t}^{\theta}=\cdot]^{1/2}$, confirming the presence of path dependence in $\sigma^{\theta}$. Moreover, the volatility level typically increases as $X_{t}^{\theta}$, in line with the leverage effect. Note also that $\hat{\sigma}^{\theta}$ is a smooth approximation of the local volatility $\sigma_{\text{loc}}$ since the neural occupation volatility has been calibrated to vanilla options.
Figure 11: Neural volatility $\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ versus spot $X_{t}^{\theta}$ for $2^{10}$ trajectories.
@@ 6 Forward Occupation Models
where $\text{OTM}(K,T)$ is the initial price of the corresponding out-of-the-money vanilla contract (put if $K<S_{t}$ and call otherwise). As $\mathfrak{o}_{0}^{T}=\partial_{T}\operatorname{\mathcal{O}}_{0}^{T}$, the instantaneous forward occupation surface can be extracted from the weighted sum of calendar spreads $\mathfrak{o}_{0}^{T}(A)=2\int_{A}\partial_{T}\text{OTM}(K,T)\frac{dK}{K^{2}}.$ An illustration is given in Fig. 13 for SPX options. In conclusion, one can can add granularity to forward variance models by including all the forward occupation times $\operatorname{\mathcal{O}}_{o}^{\cdot}(A)$, which are tradeable quantities as well.
@@ 6.1 Dynamics of Forward Occupation Flows
$\eta^{T}_{t}(A\cup A^{\prime})\cdot dW_{t}=d\mathfrak{o}_{t}^{T}(A\cup A^{\prime})=d\mathfrak{o}_{t}^{T}(A)+d\mathfrak{o}_{t}^{T}(A^{\prime})=(\eta^{T}_{t}(A)+\eta^{T}_{t}(A^{\prime}))\cdot dW_{t}.$
@@
with volatility $H_{t}^{T}=\int_{t}^{T}\eta_{t}^{s}ds=\int_{t}^{T}\eta(t,s,\mathfrak{o}_{t}^{\tau(t,s)})ds\in\operatorname{\mathcal{M}}^{d}$. Unlike $\operatorname{\mathcal{O}}$, the forward occupation flow $\operatorname{\mathcal{O}}^{T}$ does not play the role of time since its total mass process, the variance swap rate $\text{VS}^{T}$, is of infinite variation. Moreover, by leveraging the tools in Section 3, we obtain the following chain rule for functionals of $\operatorname{\mathcal{O}}^{T}$,
@@
$\varphi(\operatorname{\mathcal{O}}_{T})=(\operatorname{\mathcal{O}}_{T}(A^{\pm})-K)^{+},\quad A^{-}=(-\infty,X_{0}],\quad A^{+}=(X_{0},\infty).$